find three consecutive odd integers such that twice the sum of the first and third more than the second by fifteen
Answer by stanbon(75887)

find three consecutive odd integers such that twice the sum of the first and third more than the second by fifteen
----------------
1st: 2x-1
2nd: 2x+1
3rd: 2x+3
------------------
Equation:
2x+1 = 2(4x+2) + 15
----
2x+1 = 8x+4 + 15
2x + 1 = 8x + 19
6x = -18
x = -3
----
1st: 2x-1 = -6-1 = -7
2nd: 2x+1 = 2(-3)+1 = -5
3rd: 2x+3 = -3
-----------------
Cheers,
Stan H.
--------------